OpenStudy (anonymous):

lim -> pi /4, sin(x-pi/4)/x-pi/4

OpenStudy (freckles):

\[\lim_{x \rightarrow \frac{\pi}{4}}\frac{\sin(x-\frac{\pi}{4})}{x-\frac{\pi}{4}} \]

OpenStudy (freckles):

this reminds me of \[\lim_{u \rightarrow 0}\frac{\sin(u)}{u}=1\]

OpenStudy (anonymous):

I was thinking of that oh my goodness thank you so much!

OpenStudy (freckles):

\[\text{ if } u=x-\frac{\pi}{4} \text{ then } \lim_{x \rightarrow \frac{\pi}{4}}u=0\] so you can do a substitution if you need

OpenStudy (freckles):

you would have u->0 since x->pi/4 and u=x-pi/4

OpenStudy (freckles):

so you can put it in that exact form if you wanted

OpenStudy (anonymous):

thank you!

OpenStudy (freckles):

np